# Cycle-Consistent Bridge Diffusion Model for Accelerated MRI Reconstruction

### v1

@@ 4.5 Ablation Study
Table 4 shows an ablation study to verify the effectiveness of our designs. The traditional BBDM [25] model is set as the baseline. Results indicate that BBDM achieves better results than other diffusion models. Then, both the cycle-consistency and CDEM designs improve the performance of the knee MRI reconstruction. Meanwhile, higher SSIM results indicate better structural fidelity of the reconstruction results. Table 5 presents a comparative analysis of the settings for $t1$ and $t2$ during the training process of our CBDM. Similarly, Table 6 compares the effects of different settings for $\sqrt{\sigma_{t1}}$ and $\sqrt{\sigma_{t2}}$. The results reveal that the time steps and noise variances for the two bridge diffusion processes can vary in each step of the cyclical diffusion process, highlighting the versatility and adaptability of the CBDM training approach.
@@ 4.6 Limitations
Although the cycle-consistent design enhances the image representation, our training process still requires two bridge diffusion models for training at the same time, which may bring more training costs. Additionally, the performance in out-of-distribution experiments still drops as shown in Table 3. Therefore, future work will explore diffusion models with lower training costs and stronger generalization capabilities.
@@ 5 Conclusion
We present a cyclical bridge diffusion model for accelerated MRI reconstruction. Specifically, a cycle consistency loss is used to help preserve fine-grained details and reduce the number of diffusion time steps for the image diffusion model training. Meanwhile, the Contourlet Decomposition Embedding Module extracts structural texture information in the frequency domain, capturing image structural texture and reducing artifacts. As a result, our CBDM approach effectively restores the normal human structure and texture details of MR images. Extensive experiments demonstrate that our method achieves state-of-the-art performance to transform under-sampled MRI data to fully-sampled MRI data on the benchmark fastMRI and IXI dataset.

### v2

@@ 4.5 Ablation Study
Table 4 shows an ablation study to verify the effectiveness of our designs. The traditional BBDM [25] model is set as the baseline. Results indicate that BBDM achieves better results than other diffusion models. Then, both the self-consistency and CDEM designs improve the performance of the knee MRI reconstruction. Meanwhile, higher SSIM results indicate better structural fidelity of the reconstruction results. Table 5 presents a comparative analysis of the settings for $t1$ and $t2$ during the training process of our SC-NDB. Similarly, Table 6 compares the effects of different settings for $\sqrt{\sigma_{t2}}$ under $\sqrt{\sigma_{t1}}=1$. The results reveal that the time steps and noise variances for the two nested diffusion bridge processes can vary in each step of the diffusion bridge process, highlighting the versatility and adaptability of the SC-NDB training approach.
@@ 4.6 Limitations
Although our SC-NDB framework enhances the modeling of source image distributions, the training procedure still necessitates the simultaneous optimization of two nested diffusion bridges, potentially leading to increased computational costs. Moreover, as shown in Table 3, the performance degrades under out-of-distribution scenarios. Future work will focus on developing more computationally efficient diffusion models with improved generalization capabilities.
@@ 5 Conclusion
In this study, we proposed the Self-Consistent Nested Diffusion Bridge (SC-NDB), a novel diffusion-based framework tailored for magnitude-image-based accelerated MRI reconstruction. Unlike conventional approaches relying on complex-valued image-space or k-space data, SC-NDB focuses on the more clinically accessible magnitude images, which are widely adopted in real-world applications. By introducing a nested structure of bidirectional bridge diffusion processes and enforcing self-consistency during training, SC-NDB more effectively captures explicit priors from the source image, leading to improved translation accuracy between under-sampled and fully-sampled MRI data.
Additionally, the proposed Contour Decomposition Embedding Module (CDEM) enhances the model’s ability to retain fine anatomical structures by leveraging multi-scale and directional structural information. Extensive experiments on the fastMRI and IXI datasets demonstrate that SC-NDB consistently outperforms both state-of-the-art magnitude-based and non-magnitude-based diffusion models in terms of reconstruction quality and fidelity.
Our findings highlight the potential of SC-NDB as a practical and effective solution for real-world MRI reconstruction. Future work will explore further optimization of computational efficiency and generalization capabilities, particularly in out-of-distribution scenarios and broader clinical settings.